# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v22

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.4.0$.
@@ 1 Introduction
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [11, 12, 13, 14, 15, 16, 17]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [18, 19]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [20]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [21]. Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [22, 23]), modern natural language processing tools for the Sinhala language are scarce [24, 25]. Further, Sinhala is a diglossic language [26] which has two variations: the literary form and the spoken form.
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [24] and Wijeratne et al. [27]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [37, 38] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [39], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.4.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@ 3.2 Data Sets
XL-sum [^75] https://github.com/csebuetnlp/xl-sum/tree/master Hasan et al. [141] and M3LS [^76] https://github.com/anubhav-jangra/M3LS Verma et al. [142] are text summarizing datasets which contains Sinhala data collected from the BBC website [^77] https://www.bbc.com/sinhala. Respectively, they have 3,414 and 10,148 Sinhala text documents and their summaries. Being a multi-modal dataset, M3LS additionally contains the relevant images collected from the same source. A dynamic word-level Sinhala Sign Language video dataset of 50 classes has been published by Charuka et al. [143]. It contains 1110 videos. The OSCAR [^78] https://oscar-project.org/ dataset [144, 145] contains web mined Sinhala data consisting of 172,755,385 words in 301,066 documents.
@@
Singh et al. [148] introduced the Aya Data set [^82] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^83] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^84] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality.
@@ 3.7 Parsers
The PoS tagged data then needs to be handed over to a parser. This is an area which is not completely solved even in English due to various inherent ambiguities in natural languages. However, in the case of English, there are systems which provide adequate results [192] even if not perfect yet. The Sinhala state of affairs, is that, the first parser for the Sinhala language was proposed by Hettige and Karunananda [69] with a model for grammar [20]. The study by Liyanage et al. [46] is concentrated on the same given that they have worked on formalizing a computational grammar for Sinhala. While they do report reasonable results, yet again, do not provide any means for the public to access the data or the tools that they have developed. Kanduboda [45] have worked on Sinhala differential object markers relevant for parsing.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [111] is available to download at GitHub  [^89] http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [195] and by Senevirathne et al. [196] are not accessible to the public. Azeez and Ranathunga [197] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [111] and PoS tagging [110]Anuruddha [198] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [199], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [200] used support vector machines to detect Sinhala named entities of types person, location and, organization.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [205]. A smaller implementation of Sinhala news classification has been attempted by de Silva [24]. As mentioned in Section 3.2, their news corpus is publicly available [^90] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [206]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [207] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [208] for human interpretability of Sinhala document classification. However, they too do not provide access corpus. Nanayakkara and Ranathunga [209] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [210, 211] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [212]Demotte and Ranathunga [213] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [214]. The attempt by Sameemdeen and Selvanthan [215] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [209]Gunasekara and Haddela [210]Lakmali and Haddela [206], and Buddhika et al. [216]. Then they propose active learning [217, 218] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [219] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [29]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [220]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [221]. The study by [222] used Genetic Algorithm on Lucene [^91] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [223] uses adapter-based [224, 225, 226, 227, 228, 229] fine-tuning on XLM-R [230], for code-mixed and code-switched text. Kirindage and Godewithana [231] used LDA [232] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [233] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [101]. Both their code [^92] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9393933https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^94] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^95] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^96] https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [234] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [235] used capsule-based methods recommended by Senevirathne et al. [214] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [236] compared Word2Vec, FastText and Doc2vec [237] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [238] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [239] based on their prior work [240]. A word2vec based tool [^97] http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [241]Demotte et al. [242] proposed a sentiment analysis system based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [243], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [214] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [244] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [245] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [246] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [247]. The work by Karunarathne [248] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [249] conducted a multi model analysis on classifying Sinhala songs by emotion.
The work by Jayawickrama et al. [250, 251] used the data set released by Wijeratne and de Silva [93] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [252] and compared the results obtained with their data set against that of Senevirathne et al. [214]Aththanayaka and Naleer [253] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset [^98] https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [128]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [254] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala.

### v23

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.5.0$.
@@ 1 Introduction
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [11, 12, 13, 14, 15, 16, 17]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [18, 19]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [20]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [21]. Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [22, 23]), modern natural language processing tools for the Sinhala language are scarce [24, 25]. Further, Sinhala is a diglossic language [26] which has two variations: the literary form and the spoken form. The Vedda (Veddah) language is considered a Sinhala creole [27, 28] of a language isolate [29].
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [24] and Wijeratne et al. [30]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [40, 41] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [42], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.5.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@ 3.2 Data Sets
XL-sum [^75] https://github.com/csebuetnlp/xl-sum/tree/master Hasan et al. [144] and M3LS [^76] https://github.com/anubhav-jangra/M3LS Verma et al. [145] are text summarizing datasets which contains Sinhala data collected from the BBC website [^77] https://www.bbc.com/sinhala. Respectively, they have 3,414 and 10,148 Sinhala text documents and their summaries. Being a multi-modal dataset, M3LS additionally contains the relevant images collected from the same source. Hewapathirana et al. [146] created the first Sinhala multi-document summarizing dataset, M2DS [^78] https://huggingface.co/datasets/KushanH/m2ds which is consistent of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. They have also released the code and pre-trained models on github [^79] https://github.com/KushanMH/m2ds. A dynamic word-level Sinhala Sign Language video dataset of 50 classes has been published by Charuka et al. [147]. It contains 1110 videos. The OSCAR [^80] https://oscar-project.org/ dataset [148, 149] contains web mined Sinhala data consisting of 172,755,385 words in 301,066 documents.
@@
Singh et al. [152] introduced the Aya Data set [^84] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^85] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^86] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality. Kudugunta et al. [153] introduced MADLAD-400878787https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents. They have also released the trained checkpoints on GitHub [^88] https://github.com/google-research/google-research/tree/master/madlad_400. A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al. [154] along with data [^89] https://huggingface.co/datasets/afaji/cvqa and a leader board [^90] https://eval.ai/web/challenges/challenge-page/2305/overview.
@@ 3.7 Parsers
The PoS tagged data then needs to be handed over to a parser. This is an area which is not completely solved even in English due to various inherent ambiguities in natural languages. However, in the case of English, there are systems which provide adequate results [198] even if not perfect yet. The Sinhala state of affairs, is that, the first parser for the Sinhala language was proposed by Hettige and Karunananda [72] with a model for grammar [20]. The study by Liyanage et al. [49] is concentrated on the same given that they have worked on formalizing a computational grammar for Sinhala. While they do report reasonable results, yet again, do not provide any means for the public to access the data or the tools that they have developed. Kanduboda [48] have worked on Sinhala differential object markers relevant for parsing. A meta-study was conducted by Stephen and Zeman [199] on the Sinhala data from Universal Dependencies (UD) for identifying light verb constructions in Sinhala.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [114] is available to download at GitHub  [^95] http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [202] and by Senevirathne et al. [203] are not accessible to the public. Azeez and Ranathunga [204] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [114] and PoS tagging [113]Anuruddha [205] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [206], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [207] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [208] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [209] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [214]. A smaller implementation of Sinhala news classification has been attempted by de Silva [24]. As mentioned in Section 3.2, their news corpus is publicly available [^96] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [215]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [216] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [217] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [218] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [219, 220] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [221]Demotte and Ranathunga [222] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [223]. The attempt by Sameemdeen and Selvanthan [224] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [218]Gunasekara and Haddela [219]Lakmali and Haddela [215], and Buddhika et al. [225]. Then they propose active learning [226, 227] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [228] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [32]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [229]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [230]. The study by [231] used Genetic Algorithm on Lucene [^97] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [232] uses adapter-based [233, 234, 235, 236, 237, 238] fine-tuning on XLM-R [239], for code-mixed and code-switched text. Kirindage and Godewithana [240] used LDA [241] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [242] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [104]. Both their code [^98] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9999993https://tinyurl.com/42un7a9y, Sinhala News source classification data set100100100https://bit.ly/3PBTyfo, Sinhala News category classification data set101101101https://bit.ly/3AsfK7u, and Sinhala writing style classification data set102102102https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [243] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [244] used capsule-based methods recommended by Senevirathne et al. [223] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [245] compared Word2Vec, FastText and Doc2vec [246] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [247] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [248] based on their prior work [249]. A word2vec based tool103103103http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [250]Demotte et al. [251] proposed a sentiment analysis system based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [252], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [223] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [253] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [254] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [255] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [256]. The work by Karunarathne [257] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [258] conducted a multi model analysis on classifying Sinhala songs by emotion.
The work by Jayawickrama et al. [259, 260] used the data set released by Wijeratne and de Silva [96] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [261] and compared the results obtained with their data set against that of Senevirathne et al. [223]Aththanayaka and Naleer [262] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset104104104https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [131]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [263] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [264]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [209] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).